OpenStudy (anonymous):

Find GCF(29, 49). A) A) 1 B) B) 29 C) C) 49 D) D) 1421

OpenStudy (zzr0ck3r):

49=29*1+20 29=20*1+9 20=9*2+2 9=2*4+1 2=1*2+0 so gcf =1

OpenStudy (zzr0ck3r):

have you learned this method?

OpenStudy (zzr0ck3r):

Im sure you can see the pattern once we get to +0 you stop and take the number above it. here that is 1